Mary’s age is 3 more than twice her brother’s age. The sum of their ages is 33. What is Mary’s age?
Found 3 solutions by Alan3354, mananth, MathTherapy:
Answer by Alan3354(69443)

Mary’s age is 3 more than twice her brother’s age. The sum of their ages is 33. What is Mary’s age?
---------
M = 2B + 3
M + B = 33
---
2B + 3 + B = 33
etc

Answer by mananth(16946)

Mary’s age is 3 more than twice her brother’s age.
Let brother's age be x years
Mary's age will be (2x+3)
The sum of their ages is 33. What is Mary’s age?
x+2x+3=33
3x+3=33
subtract 3 from both sides
3x=30
divide by 3
x=10
Brother's age = 10 years
Mary's age = 2x+3 = 2(10)+3= 23 years


Answer by MathTherapy(10552)

Mary’s age is 3 more than twice her brother’s age.  The sum of their ages is 33.  What is Mary’s age?

Let Mary's and her brother's ages be M and B, respectively
   Then we get: M = 2B + 3 ---- eq (i)
      Also, M + B = 33____B = 33 - M ----- eq (ii)
                M = 2(33 - M) + 3 ----- Substituting 33 - M for B in eq (i)
                M = 66 - 2M + 3
           M + 2M = 66 + 3
               3M = 69
Mary's age, or